OpenStudy (anonymous):

dy/dx y^2=2yy'? i am pretty sure its that i just want to make sure.

OpenStudy (jdoe0001):

\(\bf \textit{using the power rule}\\ \quad \\ \cfrac{dy}{dx}\quad y^\color{red}{2}\implies \color{red}{2}y^{\color{red}{2}-1}\)

OpenStudy (anonymous):

i want the derivative with respect to x

OpenStudy (jdoe0001):

ohh

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

@jdoe0001 ?

OpenStudy (jdoe0001):

no, don't think I know that one

OpenStudy (anonymous):

oh ok

myininaya (myininaya):

\[\frac{d}{dy} y^2=2y^{2-1}\] \[\frac{d}{dx} y^2=2y^{2-1}y'\] Why do you have\[ y' y^2=2yy'\]

OpenStudy (anonymous):

i need the derivative with respect to x

myininaya (myininaya):

The derivative of what?

OpenStudy (anonymous):

y^2

myininaya (myininaya):

I already gave that. What you have is \[y'y^2=2yy'\]

OpenStudy (anonymous):

so i am right

myininaya (myininaya):

I think you mean to say dy^2/dx=2yy'

myininaya (myininaya):

There is a difference.

OpenStudy (anonymous):

ok got it thanks

OpenStudy (anonymous):

and what would it be just for y

myininaya (myininaya):

\[\frac{dy}{dx} =y'=f'(x)=\frac{df}{dx}=\frac{d}{dx} f(x)\]

myininaya (myininaya):

\[\frac{d(y^2)}{dx}=2y^{2-1} \cdot \frac{dy}{dx}=2y \frac{dy}{dx}\]